OpenStudy (anonymous):

Find the sum of a 9-term geometric sequence when the first term is 4 and the last term is 1,024 and select the correct answer below.

OpenStudy (welshfella):

first term = a1 = 4, common ratio = r 9th (last) term = a1r^8 = 4*r^8 = 1024 so 4 r^8 / 4 = 1024/4 r^8 = 1024 can you find what r is?

OpenStudy (anonymous):

is it 2.4

OpenStudy (welshfella):

sorry i made a mistake in my first post r^8 = 1024/4 = 256

OpenStudy (welshfella):

r^8 = 256 r = 2

OpenStudy (welshfella):

now we apply the formula for the sum of n terms Sn = a1 * (r^n - 1 ------ r - 1 S9 = 4 * (2^9 - 1) -------- 2 - 1

OpenStudy (welshfella):

can you work that out?

OpenStudy (welshfella):

S9 = 4 * (512 - 1)

OpenStudy (anonymous):

thank you

OpenStudy (welshfella):

yw